4. Suppose that X is a random variable having the following probability distri- bution function - 0 if r<1 1/2 if 1 x <3 1 if z 2 6 (a) Find the probability mass function of X. (b) Find the mathematical expectation and the variance of X (c) Find P(4 X < 6) and P(1 < X < 6). (d) Find E(3x -6X2) and Var(3X-4).
Answer #1

a)

below is probability mass function of X:

P(X=1)=P(X<=1)-P(X<1)=1/2-0 =1/2

P(X=3)=P(X<=3)-P(X<3)=2/3-1/2=1/6

P(X=6)=P(X<=6)-P(X<6)=1-2=1/3

b)

x P(x) xP(x) x2P(x)
1    1/2 0.500 0.500
3    1/6 0.500 1.500
6    1/3 2.000 12.000
total 3.000 14.000
E(x) =μ= ΣxP(x) = 3.0000
E(x2) = Σx2P(x) = 14.0000
Var(x)=σ2 = E(x2)-(E(x))2= 5.000

from above expectation =3

and Variance =5

c)

P(4<=X<6)=0 (as no value exist between 4 and 6 for which X has any probabilistic value)

P(1<X<6)=P(X=3)=1/6

d)

E(3X-6X2)=3E(X)-6E(X2)=3*3-6*14=-75

Var(3X-4)=9*Var(X)=9*5=45